9x+4y=-17
12y=-3-3x
using substition or elimination

Answer by ptaylor(2198)

9x+4y=-17-----------------eq1
12y=-3-3x -----------------eq2
divide each term in eq2 by 3
4y=-1-x--------------------eq2a
Now substitute 4y=-1-x from eq2a into eq1:
9x-1-x=-17 add 1 to each side
9x-1+1-x=-17+1 collect like terms
8x=-16 divide each side by 8
x=-2--substitute this into eq2a
4y=-1+2=1 divide each side by 4
y=1/4
CK
From eq1:
9*(-2)+4(1/4)=-17
-18+1=-17
-17=-17
From eq2a
4(1/4)=-1+2
1=1

Hope this helps---ptaylor